OpenStudy (anonymous):

A begger on the street can make one cigarette out of every 6 cigarette butts he finds. After one whole day of searching and checking public ashtrays the begger finds a total of 72 cigarette butts. How many cigarettes can he make and smoke from the butts he found?

OpenStudy (anonymous):

I hope no mods remove this one for no reason.

OpenStudy (jtvatsim):

Haha! Mathematics in the real-world must be the title of this textbook.

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

12

OpenStudy (bananas1234):

72 divided by 6 = 12

OpenStudy (bananas1234):

ha ha but funny question! not sure why a school ect. would ask that

OpenStudy (anonymous):

6 butts = 1 cigarette 1 butt = 1/6 cigarette 72 butts = 72 * 1/6 cigarettes = 12 cigarettes

OpenStudy (anonymous):

The answer is 14.